OpenStudy (anonymous):

does the IVP dy/dt=y*t^(1/2), y(1)=1 have a unique solution on an interval containing t=1?

OpenStudy (amistre64):

looks seperable to me

OpenStudy (amistre64):

\[ln|y| =\frac{2}{3}t^{3/2}+C \]

OpenStudy (anonymous):

IVP=initial value problem. i have y=e^(2/3t^(3/2))+1-e^(2/3) but it should be y=e^(2/3(t^(3/2)-1))...not the same

OpenStudy (amistre64):

\[\large y =exp(\frac{2}{3}t^{3/2}+C)\] \[y =exp(\frac{2}{3}t^{3/2})*exp(C)\] \[y =C_1\ exp(\frac{2}{3}t^{3/2})\] \[1 =C_1\ exp(\frac{2}{3})\] \[exp(-\frac23) =C_1\] if i see it right

OpenStudy (anonymous):

you say seperable meaning not unique? cuz it should be unique...

OpenStudy (amistre64):

seperable meaning its easy to see if theres a solution :)

OpenStudy (mathmate):

The solution is unique within an interval if the function y and its derivative are continuous on the interval.

OpenStudy (amistre64):

\[\frac{dy}{dx}=f(x)\ h(y)\] \[\frac{1}{h(y)}dy=f(x)dx\] \[\int(\frac{1}{h(y)}dy=f(x)dx)\]

OpenStudy (anonymous):

i see where i messed up. but why do you have -2/3?^^^

OpenStudy (amistre64):

divide off the e^{2/3}

OpenStudy (amistre64):

division makes an exponent go negative

OpenStudy (anonymous):

is that 1/e^(2/3)?

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

same question...\[dy/dt=6y ^{2/3}\] y(1)=0. i'm not getting 0 so i need verification. i got \[y=(2t-C)^{3}\] C=-8/19...